OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t^2 + 576t. After how many seconds does the projectile take to reach its maximum height? Show your work for full credit.

OpenStudy (irishboy123):

dh/dt = velocity = -32t +576 dh/dt = 0 at top, she it starts falling back down -32t+576 = 0 t = 576/32 = 18s check height = -16*18*18 + 576*18 = 5,185m

OpenStudy (anonymous):

Thanks!!!